OpenStudy (barrelracer17):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

do you know how to find the slope? you need that first

OpenStudy (barrelracer17):

yes i do

OpenStudy (anonymous):

what did you get?

OpenStudy (barrelracer17):

1

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

next, do you know the "point - slope" formula?

OpenStudy (barrelracer17):

its the y-y1=M(x-x1)

OpenStudy (anonymous):

right

OpenStudy (barrelracer17):

i get all of that i just think im getting my math wrong

OpenStudy (anonymous):

you can use either point \((4,1)\) or \((5,2)\) plug in the numbers

OpenStudy (anonymous):

which point did you use?

OpenStudy (barrelracer17):

for the x i used 4 and the y i used 1

OpenStudy (anonymous):

ok then \[y-1=1(x-4)\] is a first step, but really the writing the 1 is silly so you can write \[y-1=x-4\]

OpenStudy (barrelracer17):

okay

OpenStudy (anonymous):

then if you need to solve for \(y\) add \(1\) to both sides

OpenStudy (anonymous):

not sure what mistake you made before, but i hope it is clearer now

OpenStudy (barrelracer17):

i dont know either just didnt look right to me

OpenStudy (barrelracer17):

and thank you for your help

OpenStudy (anonymous):

yw